# sixth generation AE autoencoder

### v1

@@ II-F Optimization Problem
In this paper, we aim to find constellations that maximize the communication performance, i.e., the (G)MI $\mathbb{I}_{\text{(G)MI}}$ of the overall ISAC system, subject to a minimum detection probability constraint $\alpha_{\mathrm{D}}$, i.e., $P_{\mathrm{D}}\geqslant\alpha_{\mathrm{D}}$. Since all sub-carriers employ the same constellation, the total (G)MI is maximized if the (G)MI per sub-carrier is maximized, and we omit the sub-carrier index $n$ in the following. Furthermore, we reformulate the detection probability constraint $\alpha_{\mathrm{D}}$ as a kurtosis constraint $\tilde{\kappa}$, i.e., $\kappa\leqslant\tilde{\kappa}$, because the detection probability $P_{\text{D}}$ decreases monotonically with an increasing kurtosis $\kappa$ of the constellation. This leads to the following optimization problem
@@
s.t. $\textstyle\!\sum_{x}P_{\mathsf{x}}(x)=1,\quad P_{\mathsf{x}}(x)\geqslant 0,\quad\forall x\in\mathcal{X}$ (C0)
$\mathbb{E}_{\mathsf{x}}\{|\mathsf{x}|^{2}\}=1,\quad\;\;\mathbb{E}_{\mathsf{x}}\{\mathsf{x}\}=0$ (C11)
$\kappa\leqslant\tilde{\kappa}\vskip-4.26773pt$ (C22)
@@ III A Lower and Upper Bound for the maximum
$\mathbb{I}(\mathsf{x};\mathsf{y}_{\text{c}})=\mathbbm{h}(\mathsf{y}_{\text{c}})-\mathbbm{h}(\mathsf{y}_{\text{c}}|\mathsf{x})=\mathbbm{h}(\mathsf{y}_{\text{c}})-\mathbbm{h}(\mathsf{w}_{\text{c}})$ (16)
@@
$\mathbbm{h}(\mathsf{y}_{\text{c}})\geqslant\log_{2}\left(2^{\mathbbm{h}(\mathsf{x})}+2^{\mathbbm{h}(\mathsf{w}_{\text{c}})}\right),$ (17)
@@
$\mathbb{E}_{\mathsf{z}}\{|\mathsf{z}|^{4}\}=C_{2},$ (C22)
where $\mathsf{z}$ denotes either the transmit signal $\mathsf{x}$ (lower bound) or the received communication signal $\mathsf{y}_{\text{c}}$ (upper bound). The values of the corresponding moment constraints are summarized in Table I. For the upper bound, the transmit-side constraints on $\mathsf{x}$ are mapped into receive-side constraints on $\mathsf{y}_{\text{c}}$ and follow directly from substituting $\mathsf{y}_{\text{c}}=\mathsf{x}+\mathsf{w}_{\text{c}}$ into the respective moment definitions.
TABLE I: Moment constraints for entropy maximization
@@
$f_{\mathsf{z}_{\text{max}}}(z)=\mathrm{e}^{\gamma_{0}+\gamma_{2}|z|^{2}+\gamma_{4}|z|^{4}},$ (20)
@@
$C_{q}=\int_{\mathbb{C}}|z|^{2q}\mathrm{e}^{\gamma_{0}+\gamma_{2}|z|^{2}+\gamma_{4}|z|^{4}}\,\mathrm{d}z,\qquad q=0,1,2.$ (21)
We note that the optimal PDF $f_{\mathsf{z}_{\text{max}}}(z)$ is circularly symmetric with respect to the origin and, as a consequence, has zero mean, thus fulfilling constraint (C11). The resulting system of equations can be solved numerically. We show in Appendix A that $\gamma_{2}$ can be obtained by numerically solving a single nonlinear equation
@@
$\mathbbm{h}(\mathsf{z}_{\text{max}})=\frac{-\gamma_{0}-C_{1}\cdot\gamma_{2}-C_{2}\cdot\gamma_{4}}{\ln(2)},$ (25)
@@
To obtain the lower and upper bounds, the following steps must be carried out separately for each bound. First, compute the constraints $C_{0}$, $C_{2}$, and $C_{4}$ in Table I. Second, solve (22) numerically to obtain $\gamma_{2}$. Third, compute $\gamma_{0}$ (23) and $\gamma_{4}$ (24). Fourth, determine the entropy (25). Finally, insert this entropy into (18) to obtain the respective bound. We provide source code to reproduce these bounds in [27].
@@ IV Constellation Shaping using Autoencoders
In this section, we propose a bitwise AE framework that incorporates the sensing constraint (C22), enabling the joint optimization of both geometric and probabilistic shaping for S&C. The block diagram of the AE is shown in Fig. 3. In our setup, the receiver is a demapper with a Gaussian noise assumption [23], and only the constellation of the transmitter is trainable. Depending on the shaping method (geometric, probabilistic, or joint), we optimize the constellation points, their probabilities, or both. These trainable parameters, shown as orange blocks in Fig. 3, are implemented as linear layers without bias. For more details on AE-based constellation optimization, we refer the reader to [16, 15, 28].
@@
$L_{\text{sens}}={\begin{cases}0,&\kappa\leqslant\tilde{\kappa},\\ d(\kappa-\tilde{\kappa}),&\kappa>\tilde{\kappa}.\end{cases}}$ (26)
@@ V-A1 Constellation Design and Encoding Scheme
$P_{\mathsf{x}_{\text{1D}}}(x_{\text{1D}})=P_{\mathsf{x}_{\text{1D}}}(-x_{\text{1D}}),\quad\forall x_{\text{1D}}\in\mathcal{X}_{\text{1D}},\vskip-4.26773pt$ (28)
@@ V-A2 Optimizing Conventional PAS-compatible Constellations using an AE framework
Both the structural symmetry constraint and the assumption of identical marginal distributions, which reduce demapping complexity, can be integrated into our autoencoder framework. In conventional PAS, the target amplitude distribution fully determines the probability of occurrence of all constellation points. Therefore, instead of optimizing all $\tilde{M}$ probabilities $\bm{p}$, only the unconstrained parameters $\tilde{\bm{p}}_{\mathsf{a}}$, which define the target amplitude distribution $\mathrm{Softmax}(\tilde{\bm{p}}_{\mathsf{a}})=[P_{\mathsf{a}}(a_{1}),\ldots,P_{\mathsf{a}}(a_{2^{M_{\text{1D}}-1}})]$ are learned. To impose symmetry, we construct an unconstrained, symmetric 1D probability vector
$\tilde{\bm{p}}_{\mathsf{x}_{\text{1D}}}=[\tilde{\bm{p}}_{\mathsf{a}},\bm{J}_{\text{N}}\tilde{\bm{p}}_{\mathsf{a}}]\,\in\,\mathbb{R}^{2^{M_{\text{1D}}}},\vskip-4.26773pt$ (29)
@@ V-B1 Constellation Design and Encoding Scheme
using binary reflected Gray mapping $\beta_{\text{S}}:\{0,1\}^{M_{\varphi}}\rightarrow\mathcal{S}_{\text{GPAS}}$, where $\tilde{M}_{\varphi}=2^{M_{\varphi}}$ denotes the number of phase levels. Thus, each $s_{u}$ is a complex phasor with unit magnitude and quantized angle. The remaining $M_{\text{A}}=M-M_{\varphi}$ bits are used to label the amplitude using binary reflected Gray coding. This generalization transforms the conventional grid-like constellation into a circularly symmetric constellation, resembling APSK. Fig. 5 shows how different combinations of amplitude bits $M_{\text{A}}$ and phase bits $M_{\varphi}$, with a total of $M=6$ bits, influence the generalized PAS constellations and their corresponding S&C trade-off for a communication SNR of $10\text{\,}\mathrm{d}\mathrm{B}$. Among these, the configuration with $M_{\text{A}}=2$ amplitude bits and $M_{\varphi}=4$ phase bits achieves the best trade-off and is therefore selected for comparison in the following analysis. Unlike conventional PAS, generalized PAS directly produces complex-valued symbols.
@@ V-B2 Optimizing Generalized PAS-compatible Constellations using an AE framework
Similar to conventional PAS, constellations for generalized PAS can be efficiently optimized using our autoencoder framework. The probability of occurrence of the constellation points is the outer product of the trainable target amplitude distribution and a fixed uniform distribution over the phase
@@
where $\bm{1}_{\tilde{M}_{\varphi}}$ is an all-one vector of length $\tilde{M}_{\varphi}$.

### v2

@@ II-F Optimization Problem
In this paper, we aim to find constellations that maximize the communication performance, i.e., the (G)MI $\mathbb{I}_{\text{(G)MI}}$ of the overall ISAC system, subject to a minimum detection probability constraint $\alpha_{\mathrm{D}}$, i.e., $P_{\mathrm{D}}\geqslant\alpha_{\mathrm{D}}$. Since all sub-carriers employ the same constellation, the total (G)MI is maximized if the (G)MI per sub-carrier is maximized. To ease optimization, we assume that all sub-carriers observe the same noise variance $\sigma^{2}_{\text{c}}$, i.e., $\sigma^{2}_{\text{c,}n}=\sigma^{2}_{\text{c}},\ \forall n$ and we omit the sub-carrier index $n$ in the following. [^2] Frequency-selective fading results in a sub-carrier-dependent SNR, which substantially complicates the optimization and obscures the fundamental S&C trade-off that we aim to characterize. If the transmitter has no channel state information (CSI), the shaped constellation would need to generalize over a broad range of SNR values. Since the GMI-optimal constellation depends on the operating SNR, this would introduce not only the S&C trade-off studied in this work, but also an additional communications-only trade-off between low and high communication SNR operation, thereby preventing a clean characterization of S&C trade-off. Conversely, if transmitter-side CSI is available, adaptive modulation and power allocation, e.g., waterfilling, should be applied, which in turn affects the sensing performance and requires a joint optimization of constellation and power distribution. However, power allocation has only been explored recently [27] and constitutes a research topic of its own. Including power allocation in this work would significantly complicate the optimization and obscure the shaping-induced trade-off, and is therefore beyond the focus of this paper. Consequently, we assume a constant noise power across the sub-carriers. This corresponds to an AWGN communications channel and provides a clean first step for analyzing and characterizing the shaping-induced S&C trade-off, which is the focus of this work. Note that including the sub-carrier dependent SNR into the optimization is an important direction for future work.
Furthermore, we reformulate the detection-probability constraint $P_{\mathrm{D}}\geqslant\alpha_{\mathrm{D}}$ as a kurtosis constraint $\kappa\leqslant\tilde{\kappa}$. This follows from the fact that the detection probability increases monotonically with the average SINR (12), and the average SINR decreases monotonically with increasing kurtosis (21). Consequently, the detection probability $P_{\text{D}}$ is a monotonically decreasing function of the kurtosis $\kappa$ of the constellation, which directly justifies the reformulation of the detection-probability constraint as $\kappa\leqslant\tilde{\kappa}$. Although the average SINR (21), and thus the detection probability, depend on the unknown target amplitudes $|a_{j}|^{2}$, these amplitudes are not required during the constellation optimization. We exploit the monotonic relationship between kurtosis and detection probability, which holds for any given realization of the target amplitudes. In other words, decreasing the kurtosis always improves the sensing performance for any possible set of target amplitudes, so that no prior knowledge of the unknown target amplitudes is required during the optimization. [^3] The target amplitudes $|a_{j}|^{2}$ are required only to evaluate the resulting detection probability $P_{\mathrm{D}}$, but not to improve it by reducing the kurtosis $\kappa$. This leads to the following optimization problem
@@
s.t. $\textstyle\!\sum_{x}P_{\mathsf{x}}(x)=1,\quad P_{\mathsf{x}}(x)\geqslant 0,\quad\forall x\in\mathcal{X},$ (C0)
$\mathbb{E}_{\mathsf{x}}\{|\mathsf{x}|^{2}\}=1,\quad\;\;\mathbb{E}_{\mathsf{x}}\{\mathsf{x}\}=0,$ (C11)
[…]
@@ III A Lower and Upper Bound for the maximum
$\mathbb{I}(\mathsf{x};\mathsf{y}_{\text{c}})=\mathbbm{h}(\mathsf{y}_{\text{c}})-\mathbbm{h}(\mathsf{y}_{\text{c}}|\mathsf{x})=\mathbbm{h}(\mathsf{y}_{\text{c}})-\mathbbm{h}(\mathsf{w}_{\text{c}})\vskip-4.26773pt$ (23)
@@
$\mathbbm{h}(\mathsf{y}_{\text{c}})\geqslant\log_{2}\left(2^{\mathbbm{h}(\mathsf{x})}+2^{\mathbbm{h}(\mathsf{w}_{\text{c}})}\right),\vskip-2.84544pt$ (24)
@@
Here $\mathsf{x}_{\text{max}}$ and $\mathsf{y}_{\text{max,c}}$ denote the RVs associated with the entropy-maximizing PDFs.
@@
$\mathbb{E}_{\mathsf{z}}\{|\mathsf{z}|^{4}\}=C_{2},\vskip-5.69046pt$ (C22)
where $\mathsf{z}$ denotes either the transmit signal $\mathsf{x}$ (lower bound) or the received communication signal $\mathsf{y}_{\text{c}}$ (upper bound). The values of the corresponding moment constraints are summarized in Table II. For the upper bound, the transmit-side constraints on $\mathsf{x}$ are mapped into receive-side constraints on $\mathsf{y}_{\text{c}}$ and follow directly from substituting $\mathsf{y}_{\text{c}}=\mathsf{x}+\mathsf{w}_{\text{c}}$ into the respective moment definitions.
TABLE II: Moment constraints for entropy maximization
@@
$f_{\mathsf{z}_{\text{max}}}(z)=\mathrm{e}^{\gamma_{0}+\gamma_{2}|z|^{2}+\gamma_{4}|z|^{4}},\vskip-3.41418pt$ (27)
@@
$C_{q}=\int_{\mathbb{C}}|z|^{2q}\mathrm{e}^{\gamma_{0}+\gamma_{2}|z|^{2}+\gamma_{4}|z|^{4}}\,\mathrm{d}z,\qquad q=0,1,2.\vskip-2.84544pt$ (28)
We note that the optimal PDF $f_{\mathsf{z}_{\text{max}}}(z)$ is circularly symmetric with respect to the origin and, as a consequence, has zero mean, thus fulfilling constraint (C11). The resulting system of equations can be solved numerically. However, this system involves integrals over the complex plane and comprises three coupled equations in the three unknowns $\gamma_{0}$, $\gamma_{2}$, and $\gamma_{4}$, which leads to a comparatively high computational complexity. To reduce computational complexity, we reformulate the system of equations (28) in Appendix B such that $\gamma_{2}$ can be obtained by numerically solving a single nonlinear equation
@@
Consequently, determining $\gamma_{2}$ is the key step and can be achieved by solving a single nonlinear equation (29) numerically using standard one-dimensional root-finding methods.
Note that for $\kappa=2$, the optimal distribution reduces to a circular complex Gaussian, which corresponds to $\gamma_{2}=-1/C_{1}$ and $\gamma_{4}=0$. Starting from this point, we can continuously and smoothly transition from the Gaussian case ($\kappa=2$) to a unit-modulus distribution ($\kappa=1$) by gradually shifting probability mass toward the unit circle (see Fig. 14), achieved by increasing $\gamma_{2}$ and decreasing $\gamma_{4}$. This continuous transition ensures that every target value $\kappa\in(1,2]$ is attainable along this path. Moreover, if there exists a solution for the entropy-maximizing distribution, it is unique [25]. Since (29) is a reformulation of the corresponding moment-matching condition (28), the solution for $\gamma_{2}$ is likewise unique for the considered kurtosis range $\kappa\in[1,2]$, apart from potential numerical instabilities caused by vanishing or exploding terms. In such cases, the system of equations (28) can be solved instead. This does not pose a practical issue, since the bounds are computed offline and are used solely as a reference for evaluating the optimized constellations.
@@
$\mathbbm{h}(\mathsf{z}_{\text{max}})=\frac{-\gamma_{0}-C_{1}\cdot\gamma_{2}-C_{2}\cdot\gamma_{4}}{\ln(2)},\vskip-5.69046pt$ (32)
@@
To obtain the lower and upper bounds, the following steps must be carried out separately for each bound. First, compute the constraints $C_{0}$, $C_{2}$, and $C_{4}$ in Table II. Second, solve (29) numerically to obtain $\gamma_{2}$. Third, compute $\gamma_{0}$ (30) and $\gamma_{4}$ (31). Fourth, determine the entropy (32). Finally, insert this entropy into (25) to obtain the respective bound. We provide source code to reproduce these bounds in [28].
Remark: Both bounds are tight for $\kappa=2$, where they coincide with the Shannon capacity of an AWGN channel $\log_{2}(1+E_{\text{s}}/\sigma^{2}_{\text{s}})$. For $\kappa<2$, the bounds are generally loose: the lower bound is tight only when $\mathsf{x}$ is Gaussian, and the upper bound may not be attainable because the entropy-maximizing receive PDF $f_{\mathsf{y}_{\text{max,c}}}$ is not realizable as discussed in the previous remark. Nevertheless, the bounds become asymptotically tight as the communication SNR increases, since the noise power $\sigma^{2}_{\text{c}}$ decreases and the constraints $C_{1}$ and $C_{2}$ in Tab. II converge. Moreover, our numerical results in Sec. VI and Fig. 13 show that the MI of the optimized constellations approaches the upper bound closely.
@@ IV Constellation Shaping using Autoencoders
In this section, we propose to incorporate the sensing constraint (C22) into the bitwise AE framework [15], enabling the joint end-to-end optimization of both geometric and probabilistic shaping for S&C. Autoencoders have become a standard and state-of-the-art tool for constellation shaping in communication systems, see, e.g., [29, 30]. Although other machine-learning-based optimization techniques could in principle be used, we are not aware of any approach that systematically outperforms AE-based end-to-end optimization for constellation shaping. A further key advantage of the AE approach is that it naturally supports probabilistic, geometric, and joint shaping within a unified architecture, whereas many existing methods focus on only one of these aspects.
A classical AE consists of an encoder that learns an internal representation of the input and a decoder that attempts to reconstruct the input, typically both implemented as neural networks. When applied to communication systems, in particular to constellation shaping, the encoder takes the role of the transmitter by learning a representation of the binary input in the form of an optimized constellation, while the decoder takes the role of the receiver by recovering the transmitted bits. During training, encoder and decoder are optimized jointly in an end-to-end manner such that the AIR is maximized. For more details on AE-based constellation optimization, we refer the reader to [16, 15, 31].
The block diagram of the AE is shown in Fig. 4. In our setup, the receiver is a demapper with a Gaussian noise assumption [23], and only the constellation of the transmitter is trainable. Depending on the shaping method (geometric, probabilistic, or joint), we optimize the constellation points, their probabilities, or both. For geometric shaping, only the unconstrained constellation points $\tilde{\bm{x}}$ are trainable. The unconstrained probabilities $\tilde{\bm{p}}$ are initialized uniformly and remain fixed, i.e., they are not trainable. Thus, the AE optimizes the position of the constellation points while the probability of occurrence remains unchanged, i.e., uniform. For probabilistic shaping, the constellation points are initialized, e.g., as conventional $\widetilde{M}-$QAM, and remain fixed during training, while only the unconstrained probabilities $\tilde{\bm{p}}$ are trainable. This corresponds to optimizing the probability of occurrence of the constellation points while keeping their position fixed. For joint, i.e., both geometric and probabilistic, constellation shaping, the unconstrained constellation points $\tilde{\bm{x}}$ and the unconstrained probabilities $\tilde{\bm{p}}$ are trainable, which yields the most degrees of freedom. The trainable parameters, shown as orange blocks in Fig. 4, are implemented as linear layers without bias.
@@
$L_{\text{sens}}={\begin{cases}0,&\kappa\leqslant\tilde{\kappa},\\ d(\kappa-\tilde{\kappa}),&\kappa>\tilde{\kappa}.\end{cases}}\vskip-4.26773pt$ (33)
@@ V-A1 Constellation Design and Encoding Scheme
$P_{\mathsf{x}_{\text{1D}}}(x_{\text{1D}})=P_{\mathsf{x}_{\text{1D}}}(-x_{\text{1D}}),\quad\forall x_{\text{1D}}\in\mathcal{X}_{\text{1D}},$ (35)
@@ V-A2 Optimizing Conventional PAS-compatible Constellations using an AE framework
In general probabilistic shaping, the AE optimizes $\widetilde{M}$ unconstrained probabilities $\tilde{\bm{p}}$ directly (see Sec. IV and Fig. 4). However, as discussed in the previous paragraph, conventional PAS imposes a specific structure on the probability of occurrence of the constellation points: (i) the probability of a complex-valued symbol factorizes into one-dimensional probabilities for the in-phase and quadrature components, and we enforce identical marginal distributions to reduce demapping complexity, (ii) the amplitude distribution is symmetric to enforce that the sign bits are uniformly distributed. Thus, to optimize PAS-compatible constellations using the AE framework, this structure must be enforced inside the AE. Instead of learning all $\widetilde{M}$ unconstrained probabilities $\tilde{\bm{p}}$, only the unconstrained parameters $\tilde{\bm{p}}_{\mathsf{a}}$, which define the one-dimensional target amplitude distribution $\mathrm{Softmax}(\tilde{\bm{p}}_{\mathsf{a}})=[P_{\mathsf{a}}(a_{1}),\ldots,P_{\mathsf{a}}(a_{2^{M_{\text{1D}}-1}})]$, are learned. Based on this optimized target amplitude distribution, the probability of occurrence of all constellation points is then generated. To impose symmetry, we construct an unconstrained, symmetric one-dimensional probability vector
$\tilde{\bm{p}}_{\mathsf{x}_{\text{1D}}}=[\tilde{\bm{p}}_{\mathsf{a}},\bm{J}_{\text{N}}\tilde{\bm{p}}_{\mathsf{a}}]\,\in\,\mathbb{R}^{\left(2^{M_{\text{1D}}}\right)},\vskip-4.26773pt$ (36)
@@
The probability mass vector $\bm{p}_{\text{CPAS}}$ satisfies the constraints of a conventional PAS constellation and replaces the probabilities $\bm{p}$ in the AE shown in Fig. 4. This also significantly reduces the number of trainable parameters: instead of $\widetilde{M}$ unconstrained probabilities, the $2^{M_{\text{1D}}-1}$ unconstrained parameters of the one-dimensional target amplitude distribution $\tilde{\bm{p}}_{\mathsf{a}}$ define the probability of occurrence of all constellation points.
@@ V-B1 Constellation Design and Encoding Scheme
using binary reflected Gray mapping $\beta_{\text{S}}:\{0,1\}^{M_{\varphi}}\rightarrow\mathcal{S}_{\text{GPAS}}$, where $\tilde{M}_{\varphi}=2^{M_{\varphi}}$ denotes the number of phase levels. Thus, each $s_{u}$ is a complex phasor with unit magnitude and quantized angle. The remaining $M_{\text{A}}=M-M_{\varphi}$ bits are used to label the amplitude using binary reflected Gray coding. This generalization transforms the conventional grid-like constellation into a circularly symmetric constellation, resembling APSK. Fig. 6 shows how different combinations of amplitude bits $M_{\text{A}}$ and phase bits $M_{\varphi}$, with a total of $M=6$ bits, influence the generalized PAS constellations and their corresponding S&C trade-off for a communication SNR of $10\text{\,}\mathrm{d}\mathrm{B}$. The configuration with $M_{\text{A}}=2$ amplitude bits and $M_{\varphi}=4$ phase bits achieves the highest GMI across all considered kurtosis values and therefore provides the best S&C trade-off among the evaluated combinations of amplitude bits $M_{\text{A}}$ and phase bits $M_{\varphi}$. For this reason, it is selected for comparison in the following analysis. Unlike conventional PAS, generalized PAS directly produces complex-valued symbols.
@@ V-B2 Optimizing Generalized PAS-compatible Constellations using an AE framework
Similarly to conventional PAS, constellations for generalized PAS can be efficiently optimized using our AE framework by including the structural constraints of generalized PAS into the AE. The probability of occurrence of the constellation points is the outer product of the trainable target amplitude distribution and a fixed uniform distribution over the phase
@@
where $\bm{1}_{\tilde{M}_{\varphi}}$ is an all-one vector of length $\tilde{M}_{\varphi}$. The probability mass vector $\bm{p}_{\text{GPAS}}$ satisfies the constraints of a conventional PAS constellation and replaces the probabilities $\bm{p}$ in the AE shown in Fig. 4.